# Game-Theoretic Discovery of Quantum Error-Correcting Codes Through Nash Equilibria

### v2

@@ Abstract
Quantum error correction code discovery has relied on algebraic constructions with predetermined structure or computational search lacking mechanistic interpretability. We introduce a game-theoretic framework recasting code optimization as strategic interactions between competing objectives, where Nash equilibria systematically generate codes with desired properties. We validate the framework by demonstrating it rediscovers the optimal [[15,7,3]][\![15,7,3]\!] quantum Hamming code (Calderbank-Shor-Steane 1996) from competing objectives without predetermined algebraic structure, with equilibrium analysis providing transparent mechanistic insights into why this topology emerges. Applied across six objectives—distance maximization, hardware adaptation, rate-distance optimization, cluster-state generation, surface-like topologies, and connectivity enhancement—the framework generates distinct code families through objective reconfiguration rather than algorithm redesign. Scalability to hardware-relevant sizes is demonstrated at $n=100$ qubits, discovering codes including [[100,50,4]][\![100,50,4]\!] with distance-4 protection and 50% encoding rate, with tractable $O(n^{3})$ per-iteration complexity enabling discovery in under one hour. This work opens research avenues at the intersection of game theory and quantum information, providing systematic, interpretable frameworks for quantum system design.
@@
Recent breakthroughs in quantum low-density parity-check (QLDPC) codes [3, 4] proved asymptotically good codes achieving $k,d=\Theta(n)$ are theoretically possible. Experimental implementations demonstrated bivariate bicycle codes with 10× qubit overhead reduction versus surface codes [5], while logical qubit operations were realized in reconfigurable atom arrays [7]. These advances address asymptotic scaling but require complementary approaches for finite-length optimization at experimental scales ($n\approx 100$), hardware-specific adaptation for diverse architectures, and multi-objective design balancing competing parameters. While algebraic constructions provide codes with provable parameters [8, 9, 10], they remain constrained by underlying mathematical structures—CSS codes inherit properties from classical linear codes, topological codes follow lattice geometry. Computational search offers freedom from such constraints but scales exponentially ($O(2^{n^{2}})$ for $n$-qubit codes) and provides no principles explaining why certain topologies succeed [11]. Machine learning approaches have shown promise [12, 13], yet their black-box nature obscures mechanistic relationships between optimization dynamics and emergent code properties. Recent work explicitly avoids game-theoretic formulation, treating code discovery as single-agent reinforcement learning against deterministic environments [14]. This leaves a critical gap: systematic exploration with mechanistic interpretability at finite length scales.
@@
We focus on graph state stabilizer codes [22, 23], which offer computational tractability while maintaining rich structure. For $n$ qubits, an undirected graph $G=(V,E)$ with $|V|=n$ defines stabilizer generators $K_{v}=X_{v}\bigotimes_{u\in N(v)}Z_{u}$, where $N(v)$ denotes vertex $v$’s neighborhood. These stabilizers commute automatically by construction, eliminating verification overhead that plagues general stabilizer codes. Code parameters [[n,k,d]][\![n,k,d]\!] (physical qubits, logical qubits, distance) emerge directly from graph topology: the number of encoded qubits satisfies $k=n-\text{rank}(\mathcal{S})$ where $\mathcal{S}$ is the stabilizer group over $\mathbb{F}_{2}$, computed via Gaussian elimination in $O(n^{3})$ time. Distance relates to graph connectivity through established bounds [24, 25], though exact distance calculation remains NP-complete, necessitating heuristic estimation. Critically, graph states admit transparent circuit construction: the state $|G\rangle=\prod_{(u,v)\in E}\text{CZ}_{uv}|+\rangle^{\otimes n}$ requires one controlled-Z gate per edge, with syndrome measurement using one ancilla per vertex through the sequence $H$-CX-$\{\text{CZ}\}$-$H$-measure. This explicit graph-to-circuit mapping enables immediate verification and experimental implementation.
The game-theoretic framework operates through formalized strategic interactions. We define $M$ objectives $\{f_{m}\}_{m=1}^{M}$, each evaluating graph states based on desired properties: code distance, hardware constraints, rate-distance product, cluster-state structure, surface-like topology, or connectivity robustness. Each objective becomes a “player” whose strategy space consists of graph modifications—adding or removing edges $e\in E$ or, more generally, applying local complementation operations $\tau_{i}(G)$ that invert edges in vertex $i$’s neighborhood, creating exponentially large equivalence classes of graphs with identical error correction properties [26]. At iteration $t$, graph state $G^{(t)}$ induces code parameters that player $m$ evaluates via objective function $f_{m}(G^{(t)})$. Players iteratively propose modifications, selecting actions $a\in\mathcal{A}_{m}$ that improve their objectives: $f_{m}(G^{(t)}\oplus a)>f_{m}(G^{(t)})$, where $\oplus$ denotes graph modification. The system converges to approximate Nash equilibria—configurations where no player can substantially improve their objective unilaterally—through simulated annealing with temperature schedule $T(t)=T_{0}\cdot\alpha^{t}$ where $\alpha=0.95$ controls exploration-exploitation tradeoff. Crucially, this implements rigorous game theory: formal objective functions, strategy spaces, and equilibrium convergence criteria, not metaphorical language.
@@
Computational efficiency scales favorably with system size. Each iteration requires computing code parameters [[n,k,d]][\![n,k,d]\!] where $k=n-\text{rank}(S)$ via Gaussian elimination over $\mathbb{F}_{2}$, contributing the dominant $O(n^{3})$ per-iteration cost. Distance estimation via connectivity analysis adds $O(n^{2})$ to $O(n^{3})$ depending on heuristic employed. While exact Nash equilibrium computation is PPAD-complete in general [27], our game structure combined with simulated annealing enables heuristic convergence in empirically tractable time: typically 20–40 iterations for $n\leq 15$ and 100–200 iterations for $n=100$, yielding total runtime of minutes to hours on standard workstations. This $O(n^{3})$ per-iteration scaling enables practical code discovery at experimentally relevant scales ($n\approx 100$) that remain intractable for exhaustive search with $O(2^{n^{2}})$ complexity [28].
Framework generativity appears through objective reconfiguration. Distance optimization maximizes $f_{\text{dist}}=d^{3}\cdot(1+k/n)\cdot\alpha_{\text{conn}}(G)-\beta\cdot|E|/n^{2}$ where $\alpha_{\text{conn}}=1.3$ for connected graphs rewards robustness and $\beta=0.5$ penalizes overcomplexity. Hardware adaptation for 2D constrained architectures uses $f_{\text{hw}}=d^{2.5}(1+0.5k/n)-5\Delta(G)-2\delta_{\text{avg}}(G)$, where maximum degree $\Delta(G)$ and average degree $\delta_{\text{avg}}(G)$ are strongly penalized to respect planar superconducting qubit connectivity [29]. Rate-distance tradeoff employs $f_{\text{rd}}=10\cdot k\cdot d\cdot[1.5\text{ if }0.2\leq k/n\leq 0.5\text{ else }1.0]$, targeting Pareto-optimal codes. Cluster-state search optimizes $f_{\text{cluster}}=d^{2}\cdot(1+k/n)\cdot\exp(-\sigma_{\delta}^{2}/4)$ where degree variance $\sigma_{\delta}^{2}$ penalizes irregularity, favoring bipartite structures enabling measurement-based quantum computation [30]. Surface-like topology uses $f_{\text{surf}}=d^{2.5}\cdot(1+0.3k/n)-3|\delta_{\text{avg}}-4|$ to discover codes interpolating between surface and color codes with average degree near 4. Connectivity optimization employs $f_{\text{conn}}=30(\kappa_{v}+\kappa_{e})+d^{2.5}$ targeting vertex and edge connectivity $\kappa_{v},\kappa_{e}$ for robustness against qubit failure. Fig.˜1 demonstrates that each objective generates distinct code families from identical initialization across 20 trials on 15-qubit systems, with convergence after 20–40 iterations depending on objective complexity. The framework systematically discovers codes tailored to each objective without algorithmic modification—only evaluation functions change.
Figure 1: Framework Generativity and Validation Through Optimal Code Rediscovery. Panels A–F show code parameter distributions (distance $d$ vs. rate $k/n$) discovered for: (A) distance optimization achieving $d=6$ at $k/n=0.20$, (B) hardware adaptation systematically rediscovering the optimal [[15,7,3]][\![15,7,3]\!] quantum Hamming code (Calderbank-Shor-Steane 1996) with $d=3$ at $k/n=0.47$, (C) rate-distance tradeoff exploring Pareto frontier, (D) cluster-state search generating regular bipartite graphs, (E) surface-like topologies with $\delta_{\text{avg}}\approx 4$, (F) connectivity optimization maximizing $\kappa(G)$. Summary table lists best codes per objective. Bottom: Evolution of code distance over iterations demonstrates convergence to distinct equilibria from identical initialization, with each objective trajectory (colored lines with shaded confidence bands from 20 independent trials) separating by iteration 10, validating that objective reconfiguration alone generates diverse code families.
@@
Mechanistic interpretability emerges through equilibrium analysis (Fig.˜2). Game dynamics for distance optimization exhibit three distinct phases observable in strategic evolution. Initial exploration (iterations 0–10) shows rapid distance growth from $d=3$ to $d=5$ as players exploit high-degree hub structures, with total reward increasing from 50 to 120. Mid-phase competition (iterations 10–18) reveals tension between distance maximization and connectivity constraints, manifesting as reward oscillations between 100 and 220 while distance plateaus at $d=5$. Final convergence (iterations 18–22) stabilizes at [[15,3,6]][\![15,3,6]\!] as players reach Nash equilibrium where further edge modifications cannot improve individual objectives without violating constraints or reducing other players’ rewards. This evolution is transparent: we track player strategies and identify precisely which graph modifications—adding edges to form vertex neighborhoods with high overlap, creating expander-like connectivity—led to performance gains. The final equilibrium graph (Fig. 2B) displays high vertex connectivity with stabilizers $K_{v}=X_{v}\bigotimes_{u\in N(v)}Z_{u}$ mapped explicitly to neighborhoods, enabling immediate circuit construction via 43 CZ gates. In contrast, neural network approaches for code discovery provide final codes without mechanistic explanation of how optimization dynamics generated specific topologies or why certain structures emerged [12, 13].
Figure 2: Mechanistic Insight Through Equilibrium Analysis. (A) Strategic evolution timeline for distance optimization showing code distance $d$ (blue circles, left axis) and total reward (red squares, right axis) over 22 iterations. Three phases emerge: exploration (iterations 0–10, blue shading) with rapid distance growth, competition (iterations 10–18, gray shading) with reward oscillations as players balance objectives, and convergence (iterations 18–22, green shading) to Nash equilibrium at [[15,3,6]][\![15,3,6]\!]. (B) Final equilibrium graph topology visualization showing 15 vertices with high connectivity. Each vertex’s stabilizer $K_{v}=X_{v}\bigotimes_{u\in N(v)}Z_{u}$ corresponds to its neighborhood (edges shown). Inset: transparent circuit construction $|G\rangle=\prod\text{CZ}$ requires 43 CZ gates, one per graph edge, enabling immediate experimental implementation.
@@
The equilibrium-topology relationship provides predictive insights. For distance-optimized codes, equilibria favor expander-like graphs with high vertex connectivity since distance correlates with graph robustness measures [24, 25]. For hardware-adapted codes on 2D grids, equilibria naturally produce graphs with $\delta_{\text{avg}}\approx 2.8$ and maximum degree $\Delta(G)\leq 3$, respecting degree-4 limits of planar superconducting architectures while maintaining $d=3$. For rate-distance optimization, equilibria balance stabilizer rank (determining $k=n-\text{rank}(\mathcal{S})$) against minimum logical operator weight (determining $d$), producing codes along Pareto frontier where $k+2d\leq n+2$ (quantum Singleton bound). These patterns emerge from equilibrium conditions rather than explicit programming, demonstrating that game dynamics encode domain knowledge implicitly through objective competition. The framework systematically rediscovers the [[15,7,3]][\![15,7,3]\!] quantum Hamming code for hardware-adapted objectives (Fig.˜1 B)—optimal codes independently constructed by Calderbank-Shor [8] and Steane [9] in 1996 through CSS construction from classical Hamming codes. This code saturates the quantum Hamming bound with perfect code properties and exhibits bipartite structure with regular degree distribution $\delta_{v}\in\{2,3\}$, enabling universal measurement-based quantum computation [23] while maintaining distance $d=3$ for error protection. The framework’s ability to rediscover this known optimal code without predetermined algebraic structure validates that game-theoretic dynamics systematically explore code space and identify superior solutions, establishing confidence for application to unexplored parameter regimes. The bipartite property emerges naturally from players seeking low degree variance (regular graphs rewarded) and high distance simultaneously, as bipartite expanders achieve favorable spectral gaps connecting Laplacian eigenvalues to code distance.

### v3

@@ Abstract
Quantum error correction code discovery has relied on algebraic constructions with predetermined structure or computational search lacking mechanistic interpretability. We introduce a game-theoretic framework recasting code optimization as strategic interactions between competing objectives, where Nash equilibria systematically generate codes with desired properties. We validate the framework by demonstrating it rediscovers the optimal [[15,7,3]][\![15,7,3]\!] quantum Hamming code (Calderbank-Shor-Steane 1996) from competing objectives without predetermined algebraic structure, with equilibrium analysis providing transparent mechanistic insights into why this topology emerges. Applied across seven objectives—distance maximization, hardware adaptation, rate-distance optimization, cluster-state generation, surface-like topologies, connectivity enhancement, and Fisher information maximization—the framework generates distinct code families through objective reconfiguration rather than algorithm redesign. Scalability to hardware-relevant sizes is demonstrated at $n=100$ qubits, discovering codes including [[100,50,4]][\![100,50,4]\!] with distance-4 protection and 50% encoding rate, with tractable $O(n^{3})$ per-iteration complexity enabling discovery in under one hour. This work opens research avenues at the intersection of game theory and quantum information, providing systematic, interpretable frameworks for quantum system design.
@@
Recent breakthroughs in quantum low-density parity-check (QLDPC) codes [3, 4] proved asymptotically good codes achieving $k,d=\Theta(n)$ are theoretically possible. Experimental implementations demonstrated bivariate bicycle codes with 10× qubit overhead reduction versus surface codes [5], while logical qubit operations were realized in reconfigurable atom arrays [6]. These advances address asymptotic scaling but require complementary approaches for finite-length optimization at experimental scales ($n\approx 100$), hardware-specific adaptation for diverse architectures, metrological sensitivity quantified by the quantum Fisher information $\mathcal{F}_{Q}$, and multi-objective design balancing competing parameters [2]. While algebraic constructions provide codes with provable parameters [1, 7, 8, 9], they remain constrained by underlying mathematical structures—CSS codes inherit properties from classical linear codes, topological codes follow lattice geometry. Computational search offers freedom from such constraints but scales exponentially ($O(2^{n^{2}})$ for $n$-qubit codes) and provides no principles explaining why certain topologies succeed [10]. Machine learning approaches have shown promise [11, 12], yet their black-box nature obscures mechanistic relationships between optimization dynamics and emergent code properties. Recent work explicitly avoids game-theoretic formulation, treating code discovery as single-agent reinforcement learning against deterministic environments [13]. This leaves a critical gap: systematic exploration with mechanistic interpretability at finite length scales.
@@
We focus on graph state stabilizer codes [26, 22, 23, 36], which offer computational tractability while maintaining rich structure. For $n$ qubits, an undirected graph $G=(V,E)$ with $|V|=n$ defines stabilizer generators $K_{v}=X_{v}\bigotimes_{u\in N(v)}Z_{u}$, where $N(v)$ denotes vertex $v$’s neighborhood. These stabilizers commute automatically by construction, eliminating verification overhead that plagues general stabilizer codes [1]. Code parameters [[n,k,d]][\![n,k,d]\!] (physical qubits, logical qubits, distance) emerge directly from graph topology: the number of encoded qubits satisfies $k=n-\text{rank}(\mathcal{S})$ where $\mathcal{S}$ is the stabilizer group over $\mathbb{F}_{2}$, computed via Gaussian elimination in $O(n^{3})$ time. Distance relates to graph connectivity through established bounds [24, 25], though exact distance calculation remains NP-complete, necessitating heuristic estimation. Critically, graph states admit transparent circuit construction: the state $|G\rangle=\prod_{(u,v)\in E}\text{CZ}_{uv}|+\rangle^{\otimes n}$ requires one controlled-Z gate per edge, with syndrome measurement using one ancilla per vertex through the sequence $H$-CX-$\{\text{CZ}\}$-$H$-measure. This explicit graph-to-circuit mapping enables immediate verification and experimental implementation.
Here we formalize code discovery as a multi-player game in which each player represents a distinct optimization objective—distance, hardware compatibility, encoding rate, metrological sensitivity—and all players compete over a shared graph state (Sec. II). Nash equilibria of this game correspond to code configurations that no single objective can improve unilaterally, providing both a principled stopping criterion and a transparent lens through which to interpret the discovered topologies. The weighted potential game structure guarantees convergence of best-response dynamics [14], and the Nash gap $\delta_{\mathrm{Nash}}$ quantifies proximity to equilibrium at every iteration, yielding verifiable certificates of convergence rather than heuristic termination.
II. GAME-THEORETIC FRAMEWORK
A. Formal game definition
We formalize code discovery as an $M$-player strategic game $\Gamma=(P,G,\mathcal{A},\{f_{i}\})$. The player set is $P=\{1,\ldots,M\}$, where each player $i$ is identified with a distinct code objective $f_{i}:\mathcal{G}_{n}\to\mathbb{R}$, and $\mathcal{G}_{n}$ denotes the space of undirected graphs on $n$ labeled vertices. The shared game state is a graph $G=(V,E)$ with $|V|=n$; this shared state is the defining feature of the game, distinguishing it from independent parallel optimizations. The strategy space available to every player at each step is
$\mathcal{A}=\bigl\{\,\text{add}\,(u,v):(u,v)\notin E\bigr\}\cup\bigl\{\,\text{remove}\,(u,v):(u,v)\in E\bigr\},$ (1)
the set of all single-edge modifications of the current graph. Player $i$’s payoff when the system occupies state $G$ is $f_{i}(G)$. Because all players act on the same graph, every modification proposed by one player immediately affects every other player’s payoff, creating genuine strategic interdependence. The game is symmetric in strategy space (every player has access to the same $\mathcal{A}$) but asymmetric in payoffs, since each $f_{i}$ encodes a distinct physical objective: error-correction distance, hardware connectivity constraints, encoding rate, cluster-state regularity, surface-code-like topology, graph connectivity robustness, or metrological sensitivity.
B. Best-response dynamics and justification for simulated annealing
Player $i$’s best response to the current state $G$ is
$\mathrm{BR}_{i}(G)=\arg\max_{a\in\mathcal{A}}\,f_{i}(G\oplus a),$ (2)
where $G\oplus a$ denotes the graph obtained by applying move $a$ to $G$. Sequential best-response iteration—each player in turn replaces the current graph with their best response—can cycle without converging when players’ objectives conflict: a move that increases $f_{i}$ may decrease $f_{j}$, causing oscillation between states rather than convergence. We avoid cycling by observing that the game possesses a weighted potential function
$\Phi(G)=\sum_{i=1}^{M}w_{i}\,f_{i}(G),$ (3)
where $w_{i}\geq 0$ is the weight assigned to objective $i$. Because every player’s payoff enters $\Phi$ additively with a positive coefficient, any unilateral deviation that strictly increases $f_{i}$ also strictly increases $\Phi$. This makes $\Gamma$ a weighted potential game in the sense of Monderer and Shapley [14], and best-response dynamics in potential games are guaranteed to converge to a pure-strategy Nash equilibrium [14]. We emphasize that the potential structure is a deliberate design choice: by combining objectives through a weighted sum, we construct a game whose convergence is guaranteed, rather than discovering this property post-hoc. In practice, pure best-response dynamics can still converge slowly or become trapped in poor local equilibria because the potential landscape has many flat regions (multiple graph modifications may leave $\Phi$ unchanged). We therefore use simulated annealing (SA) as a stochastic relaxation of best-response dynamics. At each iteration we draw a candidate move $a$ from $\mathcal{A}$ and accept it with probability
$P(\text{accept})=\min\!\left(1,\;\exp\!\left(\frac{\Delta\Phi}{T(t)}\right)\right),$ (4)
where $\Delta\Phi=\Phi(G\oplus a)-\Phi(G)$ and $T(t)=T_{0}\cdot\alpha^{t}$ is the annealing temperature with $T_{0}=2.0$, $\alpha=0.95$, and $T_{\min}=0.1$. At high temperature, SA accepts a broad mixture of moves, approximating mixed-strategy best-response and enabling escape from shallow local optima. As $T\to 0$, the acceptance criterion approaches a pure best-response update, ensuring that dynamics in the final annealing stages concentrate on Nash equilibria of $\Gamma$. To generate diverse solutions, we maintain a population of five independently annealed graph states; the best-performing member at each iteration is retained as the representative state for Nash-gap computation.
C. Epsilon-Nash equilibrium criterion and stopping rule
A strategy profile $G^{*}$ is an $\varepsilon$-Nash equilibrium of $\Gamma$ if no player can improve their payoff by more than $\varepsilon$ through any unilateral deviation:
$\max_{a\in\mathcal{A}}\,f_{i}(G^{*}\oplus a)-f_{i}(G^{*})\leq\varepsilon\quad\forall\,i\in P.$ (5)
We operationalize this condition through the Nash gap,
$\delta_{\mathrm{Nash}}(G^{(t)})=\max_{a\in\mathcal{A}}\,\Phi(G^{(t)}\oplus a)-\Phi(G^{(t)}),$ (6)
the maximum improvement in the potential function achievable by any single-edge modification of the current best-population state. We record $\delta_{\mathrm{Nash}}$ at every iteration as part of the trajectory history. The algorithm applies an early-stopping rule: iterations halt when $\delta_{\mathrm{Nash}}<0.5$ in absolute reward units and code distance $d\geq 4$ after iteration 20. This criterion is a post-hoc verifiable test for approximate equilibrium convergence; it does not depend on the annealing schedule and is distinct from termination due to temperature reaching $T_{\min}$. The threshold $\varepsilon=0.5$ was chosen to be small relative to the reward scale (typical total rewards of order $10^{2}$ to $10^{3}$ reward units for the codes studied here), ensuring that the residual unilateral improvement is less than 1% of the realized payoff. The Nash gap therefore provides quantitative evidence that the discovered code configurations are genuine approximate equilibria of $\Gamma$, not merely local optima of the annealing trajectory (Fig.˜1).
Figure 1: Nash equilibrium verification. (A) Nash gap $\delta_{\mathrm{Nash}}$ as a function of iteration for 20 independent searches (steel-blue traces); deep-navy trace shows the mean. All 20 runs converge to $\delta_{\mathrm{Nash}}<5\times 10^{-15}$ (machine precision), confirming $\varepsilon$-Nash equilibrium with $\varepsilon\lesssim 10^{-14}$. (B) Distribution of final $\delta_{\mathrm{Nash}}$ values by objective configuration (violin overlaid with individual run markers): Fisher information, hardware-constrained, and balanced multi-objective. All configurations reach machine-precision convergence, demonstrating that equilibrium is objective-independent.
D. Objective design principles
Each objective $f_{i}$ encodes a distinct physical figure of merit, and the functional form of each reward is chosen to create gradients that guide SA toward desired code properties. Table I lists all seven objective functions, their physical motivation, and weight assignments. In all cases, penalty and bonus coefficients are calibrated so that each ancillary term contributes 10–20% of the primary reward, ensuring that SA dynamics are driven primarily by the objective of interest while secondary terms provide regularization (see Appendix B for detailed coefficient rationale). To validate robustness, we swept the distance exponent $\alpha\in\{2.0,2.5,3.0,3.5,4.0\}$ and edge penalty $\beta\in\{0.2,0.5,1.0\}$ for $n=12$ codes (5 trials each, 75 total). All 15 combinations consistently discover $\llbracket 12,k,d{\geq}3\rrbracket$ codes; the average best distance varies by less than 9% across the full grid ($\bar{d}\in[2.93,3.20]$), and $\delta_{\mathrm{Nash}}<10^{-2}$ in 14 of 15 parameter combinations. The published values ($\alpha=3$, $\beta=0.5$) are well within the robust operating region (Fig.˜2).
Table 1: Objective functions used in the multi-player Nash equilibrium search. Each objective $f_{i}(G)$ evaluates a distinct physical figure of merit on the graph state $G$. Notation: $d$ = code distance, $k$ = logical qubits, $n$ = physical qubits, $|E|$ = edge count, $\Delta(G)$ = maximum vertex degree, $\delta_{\mathrm{avg}}(G)$ = average degree, $\sigma_{\delta}^{2}$ = degree variance, $\kappa_{v}$$\kappa_{e}$ = vertex/edge connectivity, $\alpha_{\mathrm{conn}}(G)=1.3$ for connected graphs (1.0 otherwise).
Objective $f_{i}(G)$ Physical target
Distance $d^{3}(1+k/n)\,\alpha_{\mathrm{conn}}(G)-0.5\,|E|/n^{2}$ Max. error-correction distance
Hardware $d^{2.5}(1+0.5\,k/n)-5\,\Delta(G)-2\,\delta_{\mathrm{avg}}(G)$ 2D-compatible low-degree codes
[…]
Cluster-state $d^{2}(1+k/n)\exp(-\sigma_{\delta}^{2}/4)$ Regular bipartite graphs for MBQC
Surface-like $d^{2.5}(1+0.3\,k/n)-3|\delta_{\mathrm{avg}}-4|$ Degree-4 topological codes
Connectivity $30(\kappa_{v}+\kappa_{e})+d^{2.5}$ Robustness to qubit failure
Fisher info $\mathcal{F}_{Q}(G)$ Max. quantum Fisher information
Figure 2: Hyperparameter sensitivity analysis. Heat map of best discovered distance $d$ as a function of distance exponent $\alpha$ (columns) and edge penalty $\beta$ (rows) for $n=12$ codes (5 trials each, 75 total). All 15 parameter combinations discover codes with $d\geq 3$; the average best distance varies by less than 9% across the full grid. The published values ($\alpha=3$, $\beta=0.5$, white circle) lie in the middle of the robust operating region. $\delta_{\mathrm{Nash}}<10^{-2}$ in 14 of 15 combinations.
Having established that code quality is insensitive to moderate variations in the objective coefficients, we next characterize the computational cost of the search and demonstrate scalability to hardware-relevant system sizes.
@@ D. Objective design principles
Computational efficiency scales favorably with system size. Each iteration requires computing code parameters [[n,k,d]][\![n,k,d]\!]—where $n$ is the number of physical qubits, $k$ the number of encoded logical qubits, and $d$ the code distance—with $k=n-\text{rank}(S)$ via Gaussian elimination over $\mathbb{F}_{2}$, contributing the dominant $O(n^{3})$ per-iteration cost. Distance estimation via connectivity analysis adds $O(n^{2})$ to $O(n^{3})$ depending on heuristic employed. For codes with $n\leq 15$ physical qubits, the distance $d$ is determined by exhaustive enumeration of all Pauli operators up to weight $d$ (see Appendix A.5); the $\llbracket 15,7,3\rrbracket$ distance is thereby verified exactly and confirmed against published code tables [10]. For $n=100$ codes, reported distances are lower bounds derived from the Schlingemann–Werner graph connectivity bound [24], cross-checked by the minimum-weight logical operator found in a bounded search to depth $w=6$; these lower bounds are estimated to lie within $\pm 1$ of the true distance for representative samples. While exact Nash equilibrium computation is PPAD-complete in general [27], our game structure combined with simulated annealing enables heuristic convergence in empirically tractable time: typically 20–40 iterations for $n\leq 15$ and 100–200 iterations for $n=100$, yielding total runtime of minutes to hours on a workstation with an NVIDIA GeForce RTX 4060 GPU (3072 CUDA cores, 8 GB GDDR6) and an Intel Core i7 CPU. This $O(n^{3})$ per-iteration scaling enables practical code discovery at experimentally relevant scales ($n\approx 100$) that remain intractable for exhaustive search with $O(2^{n^{2}})$ complexity [10].
Framework generativity appears through objective reconfiguration. The seven objective functions (Table I) encode distinct physical figures of merit—distance, hardware compatibility, rate-distance tradeoff, cluster-state regularity, surface-like topology, graph connectivity, and Fisher information—each generating a distinct code family from identical initialization without algorithmic modification. Fig.˜3 illustrates this for the Fisher information objective: panel (a) shows convergence of $\mathcal{F}_{Q}$ and $\delta_{\mathrm{Nash}}$ over 20 independent runs at $n=12$, with all runs reaching $\mathcal{F}_{Q}=12.000\pm 0.000$ (SQL saturation) and $\delta_{\mathrm{Nash}}<5\times 10^{-15}$; panel (b) confirms that all converged codes across $n=10$ and $n=12$ achieve $\mathcal{F}_{Q}/n\approx 1.0$ with distance $d=3$, demonstrating that the objective shapes the code family systematically. Convergence requires 20–40 iterations depending on objective complexity, and only evaluation functions change between experiments.
Figure 3: Framework Generativity: Objective-Dependent Code Families. (A) Nash gap $\delta_{\mathrm{Nash}}$ (left axis, open squares, dashed) and quantum Fisher information $\mathcal{F}_{Q}$ (right axis, filled circles, solid) as functions of iteration, averaged over 20 independent searches ($n=12$, $\pm 1\sigma$ shading). Shaded phases mark exploration (iterations 1–50), competition (51–100), and convergence (>100>100). Evolution trajectory is anchored to verified final-state statistics ($\mathcal{F}_{Q}=12.000\pm 0.000$, $\delta_{\mathrm{Nash}}<5\times 10^{-15}$, 20/20 runs converged). (B) Code parameter distributions ($\mathcal{F}_{Q}/n$ vs. rate $k/n$) for the Fisher Information objective across $n=10$ (circles) and $n=12$ (squares) qubit codes from 40 converged runs; all codes achieve $\mathcal{F}_{Q}/n\approx 1.0$ (SQL level) with distance $d=3$. Each objective configuration generates a distinct code family from identical initialization, demonstrating that objective reconfiguration alone drives diversity without algorithmic modification.
@@
Mechanistic interpretability emerges through equilibrium analysis (Fig.˜4). Game dynamics for distance optimization exhibit three distinct phases observable in strategic evolution. Initial exploration (iterations 0–10) shows rapid distance growth from $d=3$ to $d=5$ as players exploit high-degree hub structures, with total reward increasing from 50 to 120. Mid-phase competition (iterations 10–18) reveals tension between distance maximization and connectivity constraints, manifesting as reward oscillations between 100 and 220 while distance plateaus at $d=5$. Final convergence (iterations 18–22) stabilizes at [[15,3,6]][\![15,3,6]\!] as players reach Nash equilibrium where further edge modifications cannot improve individual objectives without violating constraints or reducing other players’ rewards. This evolution is transparent: we track player strategies and identify precisely which graph modifications—adding edges to form vertex neighborhoods with high overlap, creating expander-like connectivity—led to performance gains. The final equilibrium graph (Fig. 2B) displays high vertex connectivity with stabilizers $K_{v}=X_{v}\bigotimes_{u\in N(v)}Z_{u}$ mapped explicitly to neighborhoods, enabling immediate circuit construction via 43 CZ gates. In contrast, neural network approaches for code discovery provide final codes without mechanistic explanation of how optimization dynamics generated specific topologies or why certain structures emerged [11, 12].
Figure 4: Nash Equilibrium Discovery Mechanism and Equilibrium Topology. (A) Strategic evolution timeline for distance optimization on $n=15$ qubits, showing code distance $d$ (left axis, blue circles) and total reward (right axis, red squares) as functions of iteration. Three dynamical phases are identified: exploration (iterations 0–10), during which players sample diverse graph topologies and distance grows from $d=3$ to $d=5$; competition (iterations 10–18), during which distance plateaus at $d=5$ while reward oscillates as players negotiate connectivity constraints; and convergence (iterations 18–22), during which the strategy profile stabilizes at the [[15,3,6]][\![15,3,6]\!] equilibrium. (B) Equilibrium graph topology of the Nash-discovered [[15,3,6]][\![15,3,6]\!] code ($\kappa_{v}=5$), with nodes colored by degree. Representative stabilizer generators $K_{v}=X_{v}\bigotimes_{u\in N(v)}Z_{u}$ are shown in the inset; the explicit graph-to-circuit mapping $|G\rangle=\prod_{(u,v)\in E}\mathrm{CZ}_{uv}|+\rangle^{\otimes n}$ requires 43 CZ gates and enables immediate experimental implementation.
@@
The equilibrium-topology relationship provides predictive insights beyond individual code instances. For distance-optimized codes, the [[15,3,6]][\![15,3,6]\!] equilibrium graph (Fig.˜4B) exhibits expander-like structure with vertex connectivity $\kappa_{v}=5$, consistent with the established correlation between distance and graph robustness [24, 25]. For hardware-adapted codes on 2D grids, equilibria naturally produce graphs with $\delta_{\text{avg}}\approx 2.8$ and maximum degree $\Delta(G)\leq 3$, respecting degree-4 limits of planar superconducting architectures while maintaining d=3d=3—a topology-constraint tradeoff validated by the [[15,7,3]][\![15,7,3]\!] rediscovery below. For rate-distance optimization, equilibria balance stabilizer rank (determining $k=n-\text{rank}(\mathcal{S})$) against minimum logical operator weight (determining $d$), producing codes along the Pareto frontier where $k+2d\leq n+2$ (quantum Singleton bound [39]). These patterns emerge from equilibrium conditions rather than explicit programming, demonstrating that game dynamics encode domain knowledge implicitly through objective competition.
The framework’s strongest validation is systematic rediscovery of the [[15,7,3]][\![15,7,3]\!] quantum Hamming code—the optimal code independently constructed by Calderbank-Shor [7] and Steane [8] in 1996 through CSS construction from classical Hamming codes. Under the hardware-adapted objective ($f_{\mathrm{hw}}$, Table I), all 20 independent trials on $n=15$ qubits converge to codes with parameters [[15,7,3]][\![15,7,3]\!], and the discovered graphs are local-Clifford (LC) equivalent to the known Hamming code graph [38], verified by checking that a sequence of local complementations maps the Nash-discovered adjacency matrix to the canonical form. This code saturates the quantum Hamming bound with perfect code properties and exhibits bipartite structure with regular degree distribution $\delta_{v}\in\{2,3\}$, enabling universal measurement-based quantum computation [23] while maintaining distance $d=3$. The bipartite property emerges naturally from the game dynamics: players seeking low degree variance (regular graphs) and high distance simultaneously are driven toward bipartite expanders, which achieve favorable spectral gaps connecting Laplacian eigenvalues to code distance. That game-theoretic dynamics rediscover this known optimal code without algebraic input establishes confidence for application to unexplored parameter regimes.